# Sensing for Free: Learn to Localize More Sources than Antennas without Pilots

### v1

@@ Abstract
Integrated sensing and communication (ISAC) represents a key paradigm for future wireless networks. However, existing approaches often require waveform modifications, dedicated pilots, or additional overhead that complicates standards integration. We propose “sensing for free”—performing multi-source localization without pilots by reusing standard uplink data symbols, where sensing happens simultaneously with data transmission, making it directly compatible with the 3GPP 5G NR and 6G specifications. With the ever-increasing number of devices in dense 6G networks, this approach becomes particularly compelling when combined with sparse arrays, which can localize a much larger number of sources compared to uniform arrays through the enlarged virtual array. However, existing pilot-free multi-source localization algorithms for sparse arrays have numerous drawbacks. They mostly first reconstruct an extended covariance matrix and then apply subspace methods, which incur prohibitive cubic complexity while being limited to second-order statistics. Performance degrades under non-Gaussian modulated data symbols and small snapshot counts. Also, the higher-order statistics that could further enhance the localization capability remain unexploited. We address these challenges with an attention-only transformer that directly processes raw signal snapshots for grid-less end-to-end direction-of-arrival (DOA) estimation. The model efficiently captures higher-order statistics while being permutation-invariant and adaptive to varying snapshot counts. Our algorithm greatly outperforms state-of-the-art artificial intelligence (AI)-based benchmarks with over 30×30$\times$30 × reduction in parameters and runtime, and enjoys excellent generalization under practical mismatches. Applied to multi-user MIMO beam training, our algorithm can localize the uplink DOAs of multiple users simultaneously in the data transmission stage. Through angular reciprocity, estimated uplink DOAs significantly prune downlink beam sweeping candidates and enhance system throughput via sensing-assisted beam management. This work demonstrates how reusing existing data transmission for sensing can enhance both multi-source localization and beam management, two key AI-for-communication use cases in 3GPP efforts towards 6G.
Index Terms:
Multi-source localization, ISAC, sparse arrays, transformers, deep learning, beam management.
@@ I-A Background
Integrated sensing and communication (ISAC) has emerged as a key paradigm for 6G wireless networks, attracting significant interest across academia, industry, and also standardization bodies [1, 2, 3, 4, 5, 6]. For instance, ISAC is recognized as one of the six major use cases of 6G [2], and has been actively studied by the 3rd generation partnership project (3GPP), e.g., TR 22.837 in Release 19 and the ongoing discussions for Release 20 [4, 5, 6]. The basic idea is to enable the communication infrastructure itself to perform sensing functions, typically by repurposing the communication signals for radar-like tasks. Existing ISAC approaches, however, often require modifications to the waveforms or frame structures to embed sensing capabilities, which can undermine the communication performance or demand upgrades to the existing infrastructure. By contrast, we are interested in a scenario where sensing can be achieved “for free” during regular data transmissions, without using additional pilots or radio resources. In particular, we focus on the uplink of a cellular system, where a multi-antenna base station (BS) attempts to localize multiple sources by processing the unknown uplink data symbols. This type of ISAC would allow next-generation networks to gain situational awareness with no extra overhead, making it easy for practical deployment.
@@
At the same time, the demand to localize a large number of sources has become increasingly prominent, driven by the proliferation of massive internet of things (IoT) [7], unmanned aerial vehicles (UAV) communications [8], and low-altitude economy [9]. Using a conventional $M$-element uniform linear array (ULA), subspace methods such as multiple signal classification (MUSIC) can only resolve at most $M-1$ uncorrelated sources since the noise subspace must be non-trivial [10]. To overcome such a limitation, sparse linear arrays (SLAs) have been proposed, such as nested [11], co-prime [12], and minimum redundancy arrays (MRAs) [13], which sparsely deploy antennas over a large aperture to enhance the sensing capability. A properly designed $M$-antenna SLA can form a much larger virtual array, enabling the localization of $O(M^{2})$ uncorrelated sources using extended covariance matrix, far more than the number of antennas $M$ [14]. These advantages have motivated interest in considering SLAs for ISAC systems [15]. In [16], the authors found that by exploiting the virtual array for localization and physical array for communications, SLAs can achieve both enhanced sensing capability and competitive communication performance. Some recent investigations also focused on the advantages of SLAs in enhancing communication performance [17, 18, 19]. In fact, SLAs represent both established and emerging paradigms in wireless communications. In MIMO systems with antenna selection [20] or switch-based hybrid analog-digital beamforming [21], activating a subset of antennas effectively creates an SLA. Moreover, recently emerging flexible-antenna technologies, such as fluid [22] and movable [23] antenna systems, deploy antennas in flexible positions within a large aperture, and can also be naturally configured as SLAs to exploit their advantages in sensing and localization. The algorithms studied in this work are hence broadly applicable. We show the variants of SLAs in Fig. 1.
@@
In this paper, we consider ISAC with SLAs and ask the following question: Can a base station localize more sources than its number of antennas during regular data transmission, without pilots or extra resources? We investigate DL-based algorithms that provide affirmative answers to this question that are accurate, fast, and robust. We further showcase the fundamental value of such “sensing for free” capability for wireless system design, in particular beam management.
@@ I-B1 Multi-Source Localization with SLAs
I-B2 Sensing-Assisted Beam Management
@@ I-C Contributions
From a system standpoint, we demonstrate that, by leveraging unknown uplink data symbols already present in each frame, multi-source localization can be achieved at the same time with data transmission, without any dedicated overhead. Because the method leaves frame structures and waveforms untouched, it could be easily embedded into forthcoming wireless systems. Exploiting the enlarged virtual aperture of SLAs, the BS can accurately localize the DOAs for more users than physical antennas. These DOA estimates, in turn, are accurate enough to enable a significant pruning of the downlink beam sweeping candidates, remarkably reducing beam management overhead and enhancing system throughput. The results in our study motivate a rethinking of how uplink data symbols can be better harnessed to facilitate the ISAC functionality.
From an algorithmic perspective, conventional multi-source localization algorithms for SLAs rely on covariance reconstruction followed by subspace methods. They are both computationally expensive and prone to performance loss under the non-Gaussian data symbols and the limited snapshots. We circumvent subspace methods with an attention-only transformer that processes raw signal snapshots of SLAs to efficiently capture higher-order statistics for grid-less end-to-end DOA estimation, avoiding both covariance reconstruction and decomposition. Our algorithm consistently outperforms state-of-the-art benchmarks across a wide signal-to-noise-ratio (SNR) range while offering a significant reduction of over 30×30$\times$30 × in computational complexity under typical system settings. It also generalizes under various practical mismatches, and scales gracefully to large-scale arrays. Extensive simulation results and ablation studies confirm these advantages, establishing a novel paradigm for pilot-free multi-source localization and ISAC with sparse arrays.
@@ I-D Paper Organization and Notation
Notation: Throughout this work, we denote scalars by italic lowercase letters, column vectors with boldface lowercase letters, while matrices by boldface uppercase letters, respectively. For any vector $\mathbf{a}$, $[\mathbf{a}]_{1:i}$ denotes its first $i$ elements. For any matrix $\mathbf{A}$, $\mathbf{A}^{*}$, $\mathbf{A}^{T}$, $\mathbf{A}^{H}$, $[\mathbf{A}]_{:,1:i}$, and $[\mathbf{A}]_{i,j}$ are its conjugate, transpose, conjugate transpose, the first $i$ columns, and the $(i,j)$-th element, respectively. The Euclidean and Frobenius norms are written as $\lVert\mathbf{a}\rVert_{2}$ and $\lVert\mathbf{A}\rVert_{F}$, respectively, and ||$\cdot$|| ⋅ | is the cardinality of a set. The operators $\operatorname{diag}(\cdot)$, $\Re({\cdot})$, and $\Im({\cdot})$ generate a diagonal matrix and extract the real and imaginary parts, respectively. The imaginary unit is $j$. The natural logarithm is denoted by $\ln(\cdot)$.
@@ II-A System Model with Sparse MIMO Arrays
Assuming that there are $K$ narrowband and far-field sources $\{s_{k}\}_{k=1}^{K}$ with a carrier wavelength $\lambda$ impinging upon a BS with a linear array from DOAs $\bm{\theta}=\{\theta_{1},\theta_{2},\ldots,\theta_{k}\}$. We assume the DOAs are independent and identically distributed (i.i.d.) and are drawn randomly from a uniform distribution $\mathcal{U}(\theta_{\min},\theta_{\max})$, where $0\leq\theta_{\min}\leq\theta_{\max}\leq\pi$ are the minimum and maximum values of the angle range, respectively. The minimum angle separation of DOAs is defined as $\Delta_{\bm{\theta},\text{min}}\triangleq\min_{i\neq j}|\theta_{i}-\theta_{j}|$. The array is equipped with $M$ antennas with an array aperture of $(N-1)d$ where $N\geq M$. It can be either uniform or sparse [^1] While we focus on SLAs, the proposed algorithms also apply to uniform arrays. This work focuses on algorithms; see [14] for sparse array design., where the antenna positions are chosen from the integer multiples of half the wavelength $d=\frac{\lambda}{2}$. We use $\bm{\Omega}=\{\Omega_{1},\Omega_{2},\ldots,\Omega_{M}\}$ to denote the indices of antennas, where $\Omega_{m}$ is an integer representing that the position of the $m$-th antenna is $\Omega_{m}d$. In general, we configure that $1=\Omega_{1}<\Omega_{2}<\ldots<\Omega_{M}=N$. For the special case of ULAs, we have $M=N$ and $\bm{\Omega}=\{1,2,\ldots,N\}$. Fig. 2 shows a comparison of uniform and sparse arrays.
The signals are transmitted from the sources to the array for $T$ snapshots. For each snapshot $t\in\{1,2,\ldots,T\}$, the received signal at the array, i.e., $\mathbf{y}_{\bm{\Omega}}(t)\in\mathbb{C}^{M\times 1}$, is given by
@@
where $p_{k}\in\mathbb{C}$, $s_{k}(t)\in\mathbb{C}$, and $\mathbf{a}_{\bm{\Omega}}(\theta_{k})\in\mathbb{C}^{M\times 1}$ are the path loss, the source signal, and the array response vector of the $k$-th source, respectively. In addition, $\mathbf{A}_{\bm{\Omega}}(\bm{\theta})=[\mathbf{a}_{\bm{\Omega}}(\theta_{1}),\mathbf{a}_{\bm{\Omega}}(\theta_{2}),\ldots,\mathbf{a}_{\bm{\Omega}}(\theta_{K})]\text{$\in\mathbb{C}^{M\times K}$}$ is the array response matrix, $\mathbf{P}=\operatorname{diag}(p_{1},p_{2},\ldots,p_{K})\in\mathbb{C}^{K\times K}$ is the matrix form of the path loss, and $\mathbf{s}(t)=[s_{1}(t),s_{2}(t),\ldots,s_{K}(t)]^{T}\in\mathbb{C}^{K\times 1}$ is the collection of source signals. The multi-user MIMO channel of the SLA is denoted by $\mathbf{H}_{\bm{\Omega}}=\mathbf{A}_{\bm{\Omega}}(\bm{\theta})\mathbf{P}\in\mathbb{C}^{M\times K}$. The noise is given by $\mathbf{n}_{\bm{\Omega}}(t)\in\mathbb{C}^{M\times 1}$. The source signals $\mathbf{s}(t)$ and the noise $\mathbf{n}_{\bm{\Omega}}(t)$ are independent of each other, and they are both temporally uncorrelated across different snapshots. The array response vector $\mathbf{\mathbf{a}}_{\bm{\Omega}}(\theta_{k})$ is
$\mathbf{\mathbf{a}}_{\bm{\Omega}}(\theta_{k})=\left[e^{j\pi(\Omega_{1}-1)\sin(\theta_{k})},\ldots,e^{j\pi(\Omega_{M}-1)\sin(\theta_{k})}\right]^{T}.$ (2)
@@
The elements of the SLA array response vector $\mathbf{\mathbf{a}}_{\bm{\Omega}}(\theta_{k})$ are selected from those of the ULA array response vector $\mathbf{\mathbf{a}}(\theta_{k})$ with $\mathbf{\mathbf{a}}_{\bm{\Omega}}(\theta_{k})=\bm{\Gamma}_{\bm{\Omega}}\mathbf{\mathbf{a}}(\theta_{k})$, where $\mathbf{\mathbf{a}}(\theta_{k})=[1,e^{j\pi\sin(\theta_{k})},\ldots,e^{j\pi(N-1)\sin(\theta_{k})}]$ and $\bm{\Gamma}_{\bm{\Omega}}\in\mathbb{R}^{M\times N}$ is a row selection matrix whose elements are ones at the $\Omega_{m}$-th position of the $m$-th column and zeros elsewhere. Similarly, we have $\mathbf{A}_{\bm{\Omega}}(\bm{\theta})=\bm{\Gamma}_{\bm{\Omega}}\mathbf{A}(\bm{\theta})$, with $\mathbf{A}(\bm{\theta})=[\mathbf{a}(\theta_{1}),\mathbf{a}(\theta_{2}),\ldots,\mathbf{a}(\theta_{K})]\text{$\in\mathbb{C}^{N\times K}$}$ being the ULA array response matrix, and $\mathbf{y}_{\bm{\Omega}}(t)=\bm{\Gamma}_{\bm{\Omega}}\mathbf{y}(t)$, where $\mathbf{y}(t)\in\mathbb{C}^{N\times 1}$ denotes the received signals by the ULA. We make the following assumptions on the source signals and the noise:
@@
  • Source signals ${s_{k}(t)}$ are all zero-mean random processes. We consider both Gaussian and non-Gaussian signals, and assume that each $s_{k}(t)$ is independent identically distributed (i.i.d.) across snapshots. The signal $s_{k}(t)$ can either be circular complex Gaussian or practical non-Gaussian modulated symbols, such as QPSK (quadrature phase shift keying) and 16QAM (quadrature modulation), etc. We assume that the source signals are mutually independent and also uncorrelated with the noise. We consider the case where the source signals $s_{k}(t)$ are unknown during the estimation process, corresponding to unknown uplink data symbols. Hence, our multi-source DOA estimation happens simultaneously with uplink data transmission [^2] The proposed algorithms can exploit any unknown uplink transmissions for localization, not necessarily limited to unknown data symbols., and requires no additional overhead and no modifications to the frame structures or waveforms.
@@
After $T$ snapshots, we stack the received signal snapshots into the matrix form. For the full $N$-element ULA, we have $\mathbf{Y}=[\mathbf{y}_{\bm{}}(1),\mathbf{y}_{\bm{}}(2),\ldots,\mathbf{y}_{\bm{}}(T)]\in\mathbb{C}^{N\times T}$, while for the $M$-element SLA, we obtain $\mathbf{Y}_{\bm{\Omega}}=[\mathbf{y}_{\bm{\Omega}}(1),\mathbf{y}_{\bm{\Omega}}(2),\ldots,\mathbf{y}_{\bm{\Omega}}(T)]\in\mathbb{C}^{M\times T},$ where $\mathbf{Y}_{\bm{\Omega}}=\bm{\Gamma}_{\bm{\Omega}}\mathbf{Y}$ through the row selection matrix $\bm{\Gamma}_{\bm{\Omega}}\in\mathbb{R}^{M\times N}$. We next define the noiseless spatial covariance matrices (SCMs). The full $N$-element ULA has noiseless SCM $\mathbf{R}=\mathbf{A}(\bm{\theta})\mathbf{P}\mathbf{A}^{H}(\bm{\theta})\in\mathbb{C}^{N\times N}$. The corresponding noiseless SCM for the $M$-element SLA is $\mathbf{R}_{\bm{\Omega}}=\bm{\Gamma}_{\bm{\Omega}}\mathbf{R}\bm{\Gamma}_{\bm{\Omega}}^{T}\in\mathbb{C}^{M\times M}$, obtained by applying the same row selection matrix $\bm{\Gamma}_{\bm{\Omega}}$. In practice, the sample SCMs of SLA and ULA are obtained based on the noisy received signal snapshots, which are respectively given by $\tilde{\mathbf{R}}_{\bm{\Omega}}\triangleq\frac{1}{T}\mathbf{Y}_{\bm{\Omega}}\mathbf{Y}_{\bm{\Omega}}^{H}$ and $\tilde{\mathbf{R}}\triangleq\frac{1}{T}\mathbf{Y}\mathbf{Y}^{H}$.

### v2

@@ Abstract
Integrated sensing and communication (ISAC) represents a key paradigm for future wireless networks. However, existing approaches often require waveform modifications, dedicated pilots, or additional overhead that complicates standards integration. We propose “sensing for free”—performing multi-source localization without pilots by reusing random and unknown uplink data symbols, where sensing happens simultaneously with data transmission, making it directly compatible with the 3GPP 5G NR and 6G specifications. With the ever-increasing number of devices in dense 6G networks, this approach becomes particularly compelling when combined with sparse arrays, which can localize a much larger number of sources compared to uniform arrays through the enlarged virtual array. However, existing pilot-free multi-source localization algorithms for sparse arrays have numerous drawbacks. They mostly first reconstruct an extended covariance matrix and then apply subspace methods, which incur prohibitive cubic complexity while being limited to second-order statistics. Performance degrades under non-Gaussian modulated data symbols in cellular wireless networks as the higher-order statistics that could further enhance the localization capability remain unexploited. We address these challenges with an attention-only transformer that directly processes raw signal snapshots for grid-less end-to-end direction-of-arrival (DOA) estimation. The model efficiently captures higher-order statistics while being permutation-invariant and adaptive to varying numbers of snapshots. Our algorithm greatly outperforms state-of-the-art artificial intelligence (AI)-based benchmarks with over $30\times$ reduction in parameters and runtime, and enjoys excellent generalization under practical mismatches. In addition, it can effectively handle multipath propagation and mixed modulation types. Beyond localization, our algorithm can also enhance multi-user MIMO beam training through angular reciprocity. The estimated DOAs in the uplink data transmission stage can significantly prune downlink beam sweeping candidates and enhance system throughput via sensing-assisted beam management. Overall, this work demonstrates how reusing existing random data payloads for sensing can enhance both multi-source localization and beam management, two key AI-for-communication use cases in 3GPP efforts towards 6G.
@@ I-A Background
Integrated sensing and communication (ISAC) has emerged as a key paradigm for 6G wireless networks, attracting significant interest across academia, industry, and also standardization bodies [1, 2, 3, 4, 5, 6, 7]. For instance, ISAC is recognized as one of the six major use cases of 6G [3], and has been actively studied by the 3rd generation partnership project (3GPP), e.g., TR 22.837 in Release 19 and the ongoing discussions for Release 20 [5, 6, 7]. The basic idea is to enable the communication infrastructure itself to perform sensing functions, typically by repurposing the communication signals for radar-like tasks. Existing ISAC approaches, however, often require modifications to the waveforms or frame structures to embed sensing capabilities, which can undermine the communication performance or demand upgrades to the existing infrastructure. By contrast, we are interested in a scenario where sensing can be achieved “for free” during regular data transmissions, without using additional pilots or radio resources. In particular, we focus on the uplink of a cellular system, where a multi-antenna base station (BS) attempts to localize multiple sources by processing the unknown uplink data symbols. This type of ISAC would allow next-generation networks to gain situational awareness with no extra overhead, making it easy for practical deployment.
@@
At the same time, the demand to localize a large number of sources has become increasingly prominent, driven by the proliferation of massive internet of things (IoT) [8], unmanned aerial vehicles (UAV) communications [9], and low-altitude economy [10]. Using a conventional $M$-element uniform linear array (ULA), subspace methods such as multiple signal classification (MUSIC) can only resolve at most M1M-1 uncorrelated sources since the noise subspace must be non-trivial [11]. To overcome such a limitation, sparse linear arrays (SLAs) have been proposed, such as nested [12], co-prime [13], and minimum redundancy arrays (MRAs) [14], which sparsely deploy antennas over a large aperture to enhance the sensing capability. A properly designed $M$-antenna SLA can form a much larger virtual array, enabling the localization of $O(M^{2})$ uncorrelated sources using extended covariance matrix, far more than the number of antennas $M$ [15]. These advantages have motivated interest in considering SLAs for ISAC systems [16]. In [17], the authors found that by exploiting the virtual array for localization and physical array for communications, SLAs can achieve both enhanced sensing capability and competitive communication performance. Some recent investigations also focused on the advantages of SLAs in enhancing communication performance [18, 19, 20]. In fact, SLAs represent both established and emerging paradigms in wireless communications. In MIMO systems with antenna selection [21] or switch-based hybrid analog-digital beamforming [22], activating a subset of antennas effectively creates an SLA. Moreover, recently emerging flexible-antenna technologies, such as fluid [23] and movable [24] antenna systems, deploy antennas in flexible positions within a large aperture, and can also be naturally configured as SLAs to exploit their advantages in sensing and localization. The algorithms studied in this work are relevant to different practical scenarios [^1] For fluid/movable antennas, antenna positions are usually set at the start of each coherence interval and then kept fixed during data transmission [25, 26]. The algorithms operate on a block of $T$ snapshots within that interval, which is similar to static SLAs.. We show the variants of SLAs in Fig. 1.
@@
In this paper, we study multi-source localization with SLAs using unknown uplink data symbols and ask the following question: Can a base station localize more sources than its number of antennas during regular data transmission, without pilots or extra resources? We investigate DL-based algorithms that provide affirmative answers to this question that are accurate, fast, and robust. We further showcase the fundamental value of such “sensing for free” capability for wireless system design, in particular beam management.
@@ I-B1 Multi-Source Localization with SLAs
Parallel to developments from the algorithm side, recent non-asymptotic analysis also provides finite-sample guarantees related to DOA estimation with sparse arrays. In [38], Sarangi et al. established a precise trade-off between temporal and spatial measurements using high-dimensional concentration bounds, and importantly, dispelled the common belief that super-resolution with virtual arrays intrinsically demands a very large number of snapshots, which was followed by [39] that presented a non-asymptotic Cramér-Rao bound (CRB) analysis under a total-power constraint across antennas. These results show that accurate sparse array DOA estimation can be achieved with a practical number of snapshots available in cellular wireless networks.
I-B2 Sensing with Random Communication Signals
Frame structures of modern wireless communication systems devote only a small fraction of time-frequency resources to deterministic and known pilots, while most of the frame consists of random and unknown data symbols. Some recent works studied the theoretical benefits of combining those random signals with deterministic pilots for different sensing problems. In [40], the authors formulated bistatic target detection with both deterministic pilots and random payloads and derived a generalized likelihood ratio test (GLRT)-based detector with an asymptotic false-alarm rate analysis. In [41], the authors studied monostatic ISAC using both pilots and random data for sensing, and studied an ISAC precoder optimization problem. In [42], the authors derived a Ziv-Zakai bound for orthogonal frequency division multiplexing (OFDM) time-of-arrival estimation using both pilots and unknown data, and quantified the performance gains over pilot-only sensing. These theoretical studies have shown the potentials of random communication signals for enhancing pilot-based sensing.
I-B3 Sensing-Assisted Beam Management
@@ I-C Contributions
From a system standpoint, we demonstrate that, by leveraging unknown uplink data symbols already present in each frame, multi-source localization can be achieved at the same time with data transmission, without any dedicated overhead. Because the method leaves frame structures and waveforms untouched, it could be easily embedded into forthcoming wireless systems. Exploiting the enlarged virtual aperture of SLAs, the BS can accurately localize the DOAs for more users than physical antennas. These DOA estimates, in turn, are accurate enough to enable a significant pruning of the downlink beam sweeping candidates, remarkably reducing beam management overhead and enhancing system throughput. The sensing-assisted beam management is also shown to work robustly under uplink-downlink reciprocity mismatch. The results in our study motivate a rethinking of how uplink data symbols can be better harnessed to facilitate the sensing functionality.
From an algorithmic perspective, conventional multi-source localization algorithms for SLAs rely on covariance reconstruction followed by subspace methods. They are limited to second-order statistics, computationally expensive and prone to performance loss under the non-Gaussian data symbols and the limited snapshots. We circumvent subspace methods with an attention-only transformer that processes raw signal snapshots of SLAs to efficiently capture higher-order statistics for grid-less end-to-end DOA estimation, avoiding the high-complexity covariance reconstruction and decomposition steps. Our algorithm consistently outperforms state-of-the-art benchmarks across a wide signal-to-noise-ratio (SNR) range while offering a significant reduction of over $30\times$ in computational complexity under typical system settings. It also generalizes under various practical mismatches, scales gracefully to large-scale MIMO arrays, and effectively handles multipath propagation and mixed modulation types. Extensive simulation results confirm these advantages, establishing a novel paradigm for pilot-free multi-source localization with sparse arrays. Ablation studies are also provided to investigate the reasons behind the performance gains.
To avoid repetition later, we summarize the scope of this work in the following remark.
Remark 1 (“sensing for free”).
Throughout this paper, the BS opportunistically reuses the random and unknown uplink data payloads to perform multi-source localization in parallel with data transmission, without introducing any new pilots, signaling, or frame changes. The resulting uplink DOA estimates serve as side information to prune the downlink beam sweeping candidates, thereby reducing beam management overhead and improving system throughput.
@@ I-D Paper Organization and Notation
Notation: Throughout this work, we denote scalars by italic lowercase letters, column vectors with boldface lowercase letters, while matrices by boldface uppercase letters, respectively. For any vector $\mathbf{a}$, $\mathbf{a}]_{1:i}$ denotes its first $i$ elements. For any matrix $\mathbf{A}$, $\mathbf{A}^{*}$, $\mathbf{A}^{\mathsf{T}}$, $\mathbf{A}^{\mathsf{H}}$, $\mathbf{A}]_{:,1:i}$, and $\mathbf{A}]_{i,j}$ are its conjugate, transpose, conjugate transpose, the first $i$ columns, and the $(i,j)$-th element, respectively. The Euclidean and Frobenius norms are written as $\lVert\mathbf{a}\rVert_{2}$ and $\lVert\mathbf{A}\rVert_{F}$, respectively, and $|\cdot|$ is the cardinality of a set. The operators $\operatorname{diag}(\cdot)$, $\Re({\cdot})$, and $\Im({\cdot})$ generate a diagonal matrix and extract the real and imaginary parts, respectively. The imaginary unit is $j$. The natural logarithm is denoted by $\ln(\cdot)$.
@@ II-A System Model with Sparse MIMO Arrays
Assuming that there are $K$ narrowband and far-field sources $\{s_{k}\}_{k=1}^{K}$ with a carrier wavelength $\lambda$ impinging upon a BS with a linear array from DOAs $\bm{\theta}=\{\theta_{1},\theta_{2},\ldots,\theta_{K}\}$ in the uplink data transmission stage, as shown in Fig. 2. We assume the DOAs are independent and identically distributed (i.i.d.) and are drawn randomly from a uniform distribution $\mathcal{U}(\theta_{\min},\theta_{\max})$, where $0\leq\theta_{\min}\leq\theta_{\max}\leq\pi$ are the minimum and maximum values of the angle range, respectively. The DOAs also satisfy a minimum angle separation constraint defined as $\Delta_{\bm{\theta},\text{min}}\triangleq\min_{i\neq j}\left|\theta_{i}-\theta_{j}\right|,\forall i,j\in\{1,2,\ldots,K\}$. The array is equipped with $M$ antennas with an array aperture of $Nd$ where $M\leq N$. It can either be uniform or sparse [^2] While we focus on SLAs, the proposed algorithms also apply to uniform arrays. This work focuses on algorithms; see [15] for sparse array design., where the antenna positions are chosen from the integer multiples of half the wavelength $d=\frac{\lambda}{2}$. We use $\bm{\Omega}=\{\Omega_{1},\Omega_{2},\ldots,\Omega_{M}\}$ to denote the indices of antennas, where $\Omega_{m}$ is an integer representing that the position of the $m$-th antenna is $\Omega_{m}d$. In general, we configure that $1=\Omega_{1}<\Omega_{2}<\ldots<\Omega_{M}=N$. For the special case of ULAs, we have $M=N$ and $\bm{\Omega}=\{1,2,\ldots,N\}$. Fig. 3 shows the system model and a comparison of uniform and sparse arrays.
Figure 2: System model. In the uplink, there are $K$ sources impinging upon a BS from DOAs $\bm{\theta}=\{\theta_{1},\theta_{2},\ldots,\theta_{K}\}$. The BS aims to localize these sources based on the received random and unknown uplink data payloads.
The signals are transmitted from the sources to the array for $T$ snapshots. For each snapshot index $t\in\{1,2,\ldots,T\}$, the received signal at the array, i.e., $\mathbf{y}_{\bm{\Omega}}(t)\in\mathbb{C}^{M\times 1}$, is
@@
where $p_{k}\in\mathbb{C}$, $s_{k}(t)\in\mathbb{C}$, and $\mathbf{a}_{\bm{\Omega}}(\theta_{k})\in\mathbb{C}^{M\times 1}$ are the path loss, the source signal, and the array response vector of the $k$-th source, respectively. In addition, $\mathbf{A}_{\bm{\Omega}}(\bm{\theta})=[\mathbf{a}_{\bm{\Omega}}(\theta_{1}),\mathbf{a}_{\bm{\Omega}}(\theta_{2}),\ldots,\mathbf{a}_{\bm{\Omega}}(\theta_{K})]\text{$\in\mathbb{C}^{M\times K}$}$ is the array response matrix, $\mathbf{P}=\operatorname{diag}(p_{1},p_{2},\ldots,p_{K})\in\mathbb{C}^{K\times K}$ is the matrix form of the path loss, and $\mathbf{s}(t)=[s_{1}(t),s_{2}(t),\ldots,s_{K}(t)]^{\mathsf{T}}\in\mathbb{C}^{K\times 1}$ is the collection of source signals. The multi-user MIMO channel of the SLA is denoted by $\mathbf{H}_{\bm{\Omega}}=\mathbf{A}_{\bm{\Omega}}(\bm{\theta})\mathbf{P}\in\mathbb{C}^{M\times K}$. The noise is given by $\mathbf{n}_{\bm{\Omega}}(t)\in\mathbb{C}^{M\times 1}$. The source signals $\mathbf{s}(t)$ and the noise $\mathbf{n}_{\bm{\Omega}}(t)$ are independent of each other, and they are both temporally uncorrelated across different snapshots. The array response vector $\mathbf{\mathbf{a}}_{\bm{\Omega}}(\theta_{k})$ is
$\mathbf{\mathbf{a}}_{\bm{\Omega}}(\theta_{k})=\left[e^{j\pi(\Omega_{1}-1)\sin(\theta_{k})},\ldots,e^{j\pi(\Omega_{M}-1)\sin(\theta_{k})}\right]^{\mathsf{T}}.$ (2)
@@
The elements of the SLA array response vector $\mathbf{\mathbf{a}}_{\bm{\Omega}}(\theta_{k})$ are selected from those of the ULA array response vector $\mathbf{\mathbf{a}}(\theta_{k})$ with $\mathbf{\mathbf{a}}_{\bm{\Omega}}(\theta_{k})=\bm{\Gamma}_{\bm{\Omega}}\mathbf{\mathbf{a}}(\theta_{k})$, where $\mathbf{\mathbf{a}}(\theta_{k})=[1,e^{j\pi\sin(\theta_{k})},\ldots,e^{j\pi(N-1)\sin(\theta_{k})}]$ and $\bm{\Gamma}_{\bm{\Omega}}\in\mathbb{R}^{M\times N}$ is a row selection matrix whose elements are ones at the $\Omega_{m}$-th position of the $m$-th row and zeros elsewhere. Similarly, we have $\mathbf{A}_{\bm{\Omega}}(\bm{\theta})=\bm{\Gamma}_{\bm{\Omega}}\mathbf{A}(\bm{\theta})$, with $\mathbf{A}(\bm{\theta})=[\mathbf{a}(\theta_{1}),\mathbf{a}(\theta_{2}),\ldots,\mathbf{a}(\theta_{K})]\text{$\in\mathbb{C}^{N\times K}$}$ being the ULA array response matrix, and $\mathbf{y}_{\bm{\Omega}}(t)=\bm{\Gamma}_{\bm{\Omega}}\mathbf{y}(t)$, where $\mathbf{y}(t)\in\mathbb{C}^{N\times 1}$ denotes the received signals by the ULA. We make the following assumptions on the source signals and the noise:
@@
  • Source signals $s_{k}(t)$} are all zero-mean random processes. We consider both Gaussian and non-Gaussian signals, and assume that each $s_{k}(t)$ is independent identically distributed (i.i.d.) across snapshots. The signal $s_{k}(t)$ can either be circular complex Gaussian or practical non-Gaussian modulated symbols, such as QPSK (quadrature phase shift keying) and 16QAM (quadrature modulation), etc. We assume that the source signals are mutually independent [^3] The proposed algorithm in this work is applicable to not only independent sources but also coherent sources arising from multipath propagation [36]. The case of coherent sources will be discussed separately in Section V-G. and also uncorrelated with the noise, if not otherwise stated. We consider the case where the source signals $s_{k}(t)$ are unknown during the estimation process, corresponding to unknown uplink data symbols. Hence, our multi-source DOA estimation happens simultaneously with uplink data transmission, and requires no additional overhead and no modifications to the frame structures or waveforms.
@@
After $T$ snapshots [^4] We assume DOAs are unchanged within a short block of at least $T$ symbols. In the simulations, we use $T=50$ by default and also report results with value as small as $T=10$. The proposed method does not rely on long continuous traffic and remains applicable when only brief bursts exist. For very sparse or sporadic traffic, activity detection and temporal aggregation could be combined as an extension, which we leave for future work., we stack the received signal snapshots into the matrix form. For the full $N$-element ULA, we have $\mathbf{Y}=[\mathbf{y}_{\bm{}}(1),\mathbf{y}_{\bm{}}(2),\ldots,\mathbf{y}_{\bm{}}(T)]\in\mathbb{C}^{N\times T}$, while for the $M$-element SLA, we obtain $\mathbf{Y}_{\bm{\Omega}}=[\mathbf{y}_{\bm{\Omega}}(1),\mathbf{y}_{\bm{\Omega}}(2),\ldots,\mathbf{y}_{\bm{\Omega}}(T)]\in\mathbb{C}^{M\times T},$ where $\mathbf{Y}_{\bm{\Omega}}=\bm{\Gamma}_{\bm{\Omega}}\mathbf{Y}$. We next define the noiseless spatial covariance matrices (SCMs). The full $N$-element ULA has noiseless SCM $\mathbf{R}=\mathbf{A}(\bm{\theta})\mathbf{P}\mathbf{A}^{\mathsf{H}}(\bm{\theta})\in\mathbb{C}^{N\times N}$. The corresponding noiseless SCM for the $M$-element SLA is $\mathbf{R}_{\bm{\Omega}}=\bm{\Gamma}_{\bm{\Omega}}\mathbf{R}\bm{\Gamma}_{\bm{\Omega}}^{\mathsf{T}}\in\mathbb{C}^{M\times M}$. In practice, the sample SCMs of SLA and ULA are obtained based on the noisy received signal snapshots, which are respectively given by $\tilde{\mathbf{R}}_{\bm{\Omega}}\triangleq\frac{1}{T}\mathbf{Y}_{\bm{\Omega}}\mathbf{Y}_{\bm{\Omega}}^{\mathsf{H}}$ and $\tilde{\mathbf{R}}\triangleq\frac{1}{T}\mathbf{Y}\mathbf{Y}^{\mathsf{H}}$.